OpenStudy (anonymous):

Find a number that is between 7/11 and 7/10.

ganeshie8 (ganeshie8):

Hint : \[\dfrac{7}{11} = \dfrac{70}{110}\] \[\dfrac{7}{10} = \dfrac{77}{110}\]

ganeshie8 (ganeshie8):

that means you need to choose a number between \(\large \dfrac{70}{110}\) and \(\large \dfrac{77}{110}\)

OpenStudy (anonymous):

76/99, 77/99, 25/44, or 23/33